# Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA

### v1

@@ Abstract
We present GSCDB138, a rigorously curated benchmark library of 138 data sets (8383 entries) covering main-group and transition-metal reaction energies and barrier heights, non-covalent interactions, dipole moments, polarizabilities, electric-field response energies, and vibrational frequencies. Legacy data from GMTKN55 and MGCDB84 have been updated to today’s best reference values; redundant, spin-contaminated, or low-quality points were removed, and many new, property-focused sets were added. Testing 29 popular density-functional approximations shows the expected Jacob’s-ladder hierarchy overall, yet reveals interesting exceptions: r2SCAN-D4 (meta-GGA) rivals hybrids for frequencies, and electric-field errors correlate poorly with ground-state energetics. B97M-V and B97X-V are the most balanced hybrid meta-GGA and hybrid GGA, respectively; B97M-V and revPBE-D4 lead the meta-GGA and GGA classes. Double hybrids lower mean errors by about 25 % versus the best hybrids but demand careful frozen-core, basis-set, and multi-reference treatment. GSCDB138 offers a comprehensive, openly documented platform for stringent DFA validation and for training the next generation of non-empirical and machine-learned functionals.
@@ 1 Introduction
As more research groups began generating high-quality (and sometimes low-quality) benchmark data, the need to systematically integrate and curate these efforts into a larger, validated, and community-accessible database became increasingly apparent. A landmark effort was the GMTKN23 and GMTKN30 databases 37, 38, which cover general main-group thermochemistry, kinetics, and noncovalent interactions. In 2017, the larger GMTKN55 compilation was introduced 10 and is now employed to evaluate nearly all modern density functionals 10, 39, 40. That same year, the even larger Main Group Chemistry Database (MGCDB84) was also presented, providing a comprehensive benchmark of over 200 density functionals 16. Other significant efforts were also reported.41, 42
@@
Given that these state-of-the-art databases are now about eight years old, there is a significant opportunity to improve diversity and quality of data in new compilations. This work reports the result of our effort to develop a larger and more comprehensive database, Gold-Standard Chemical Database 138 (GSCDB138). GSCDB138 aims to deliver gold-standard accuracy and contains 138 datasets encompassing a total of 8,383 individual data points (requiring 14,013 single-point energy calculations). Key improvements in diversity include extensive transition-metal data drawn from realistic organometallic reactions 35, 43, 44 and well-defined model complexes 45, 46. A second important addition is energy differences that reflect molecular properties. Recently, there has been controversy regarding the accuracy of electron densities from modern functionals 47 and thus we add density-dependent properties, such as dipole moments 48, polarizabilities 49, 50, and the field-dependence of energies 51. Additionally, we introduce datasets for vibrational frequencies 52, further extending the range of available benchmarking targets. Beyond these brand-new energy categories, we also expand the coverage and diversity within established categories. Furthermore, we carefully and systematically prune all questionable data points, especially those potentially affected by spin-symmetry-breaking 53, 54, 55, yielding a benchmark suite aimed at gold-standard accuracy across an unprecedented range of chemistry.
@@
The outline of the paper is as follows. In the next section, we provide an overview of the Gold-Standard Chemical Database 138 (GSCDB138), including a comprehensive description of all 138 benchmark sets, their characteristics, and reference values (Table LABEL:chap5tab:overview). In Section 3, we discuss important considerations in constructing the database and choosing reference values. This includes the integration of existing benchmark databases (MGCDB84 and GMTKN55), updates and adjustments to reference values, resolution of spin contamination issues, inclusion of newly added data sets—particularly for transition metals and molecular properties—and basis set considerations. In Section 4, we present a DFT benchmark study that evaluates the performance of 29 selected density functionals across the new database, identifying the strengths and limitations of the functionals themselves and their corresponding functional design. Finally, in Section 5, we present our conclusions, summarizing the key findings and discussing the implications for future density functional development.
@@ 2 Overview of the Database
Table LABEL:chap5tab:overview is a simple description of all 138 benchmark sets, including a short description, the number of data points contained in each set, the root-mean-square value of their energy differences, and the reference. A more detailed version is attached in the Supplemental Information.
@@
BH28 | Highly accurate subset chosen from BHPERI, CRBH20, BHDIV10, and PX13 sets | 28 | 35.18 | 56 | |
@@
BH876 | Comprehensive reaction barrier heights | 876 | 26.93 | 57 | |
@@
BHPERI11 | Barrier heights of pericyclic reactions | 11 | 23.00 | 58 | 37 |
@@
CRBH14 | Barrier heights for cycloreversion of heterocyclic rings | 14 | 45.85 | 59 | |
DBH22 | Highly accurate subset of BH76 | 22 | 29.22 | 60 | 32 |
INV23 | Inversion/Racemization barrier heights | 23 | 36.57 | 61 | |
ORBH35 | Difficult barrier heights for oxygen reactions (e.g., vinylperoxy radical and CO3) | 35 | 32.86 | 62 | |
PX9 | Proton-exchange barriers in H2O, NH3, and HF clusters | 9 | 38.87 | 63 | |
WCPT26 | Barrier heights of water-catalyzed proton-transfer reactions | 26 | 36.11 | 64 | |
@@
Dip146 | Dipole moments for 152 small systems | 190 | 0.12 | 48 | |
HR46 | Static polarizabilities for 46 systems | 128 | 0.47 | 50 | 65 |
OEEF | Relative energies in oriented external electric fields compared to zero field | 128 | 18.07 | 51 | |
Pol130 | Static polarizabilities for 132 small systems | 296 | 1.64 | 49, 66 | |
T144 | Static polarizabilities for 144 systems | 421 | 0.69 | 50 | 67 |
V30 | Frequencies of small molecular dimers with different polarity combinations (polar-polar, polar-nonpolar, and nonpolar-nonpolar) | 275 | 6.0e-4 | 52 | |
@@
A19Rel6 | Relative energies of 19 complexes from the A24 set on potential energy curves (PECs) | 114 | 1.65 | 68 | |
ACONF | Isomerization energies of alkane conformers | 15 | 2.23 | 69 | |
AlkIsomer11 | Isomerization energies of alkanes with 4-8 carbon atoms | 11 | 1.81 | 70 | |
Amino20x4 | Isomerization energies of amino acid conformers | 80 | 2.98 | 71 | |
BUT14DIOL | Isomerization energies of butane-1,4-diol conformers | 64 | 2.91 | 10 | 72 |
C20C246 | Isomerization energies of C20 and C24 | 6 | 41.54 | 73 | |
C60ISO7 | Isomerization energies of C60 isomers | 7 | 102.22 | 74 | |
DIE60 | Isomerization energies for double-bond migration in conjugated dienes | 60 | 5.06 | 75 | |
EIE22 | Isomerization energies of enecarbonyls | 22 | 4.97 | 76 | |
H2O16Rel4 | Isomerization energies of (H2O)16 (boat and fused cube structures) | 4 | 0.45 | 77 | |
H2O20Rel9 | Isomerization energies of (H2O)20 (low-energy structures) | 9 | 2.76 | 78 | 79 |
ICONF | Isomerization energies in conformers of inorganic systems | 17 | 4.54 | 10 | |
@@
ISO34 | Isomerisation energies of small- and medium-sized organic molecules | 34 | 20.30 | 10 | 80 |
ISOL23 | Isomerisation energies of large organic molecules | 23 | 28.03 | 10 | 81 |
ISOMERI- ZATION20 | Isomerization energies from the W4-11 data set | 20 | 44.05 | 82 | |
MCONF | Isomerization energies of melatonin conformers | 51 | 5.36 | 10 | 83 |
PArel | Isomerization energies of protonated isomers | 20 | 6.23 | 10 | |
PCONF21 | Isomerization energies of tri- and tetrapeptide conformers | 18 | 1.79 | 10 | 84, 85 |
Pentane13 | Isomerization energies of stationary points on the n-pentane torsional surface | 13 | 6.75 | 86 | |
S66Rel7 | Relative energies of 66 complexes from the S66 set on PECs | 462 | 2.74 | 87 | 35 |
SCONF | Isomerization energies of sugar conformers | 17 | 4.96 | 10 | 88 |
Styrene42 | Isomerization energies of C8H8 | 42 | 68.86 | 89 | |
SW49Rel28 | Isomerization energies of SO42-(H2O)n (n = 3-5) | 28 | 1.55 | 90 | |
TAUT15 | Isomerization energies of tautomers | 15 | 4.52 | 10 | |
UPU23 | Isomerization energies of RNA-backbone conformers | 23 | 6.52 | 91 | |
@@
3B-69 | 3-Body non-additive interaction energies in three different orientations of 23 molecular crystals | 69 | 0.69 | 92 | |
3BHET | 3-Body non-additive interaction energies in molecular trimers | 20 | 2.01 | 93 | |
A24 | Highly accurate binding energies of small non-covalent complexes | 24 | 2.65 | 94 | 95 |
ADIM6 | Interaction energies of n-alkane dimers | 6 | 3.66 | 10 | 96 |
AHB21 | Interaction energies in anion-neutral dimers | 21 | 26.17 | 97 | |
Bauza30 | Binding energies of halogen-, chalcogen-, and pnicogen-bonded dimers | 30 | 23.65 | 98 | 99 |
BzDC215 | PECs for benzene interacting with two rare-gas atoms and eight first- and second-row hydrides | 215 | 1.81 | 100 | |
@@
CHB6 | Interaction energies in cation-neutral dimers | 6 | 27.85 | 97 | |
CT20 | Binding energies of charge-transfer complexes | 20 | 1.07 | 101 | |
DS14 | Binding energies of complexes containing divalent sulfur | 14 | 3.70 | 102 | |
FmH2O10 | Binding energies of isomers of F-(H2O)10 | 10 | 168.50 | 78 | 103 |
@@
HEAVY28 | Binding energies between heavy element hydrides | 28 | 1.43 | 10 | 96 |
@@
HW6Cl5 | Binding energies of Cl-(H2O)n (n = 1-6) | 5 | 62.84 | 78 | 103 |
HW6F | Binding energies of F-(H2O)n (n =1-6) | 6 | 81.42 | 78 | 103 |
@@
IL16 | Interaction energies in anion-cation dimers | 16 | 109.34 | 97 | |
@@
Shields38 | Binding energies of (H2O)n (n = 2 - 10) | 38 | 51.61 | 124, 125 | 126 |
SW49Bind22 | Binding energies of isomers of SO42-(H2O) (n = 3 -5) | 22 | 34.20 | 90 | |
TA13 | Binding energies of dimers involving radicals | 13 | 22.00 | 127 | |
WATER27 | Binding energies in (H2O)n, H+(H2O)n and OH-(H2O)n | 27 | 101.18 | 124, 125 | 128 |
X40 | Binding energies of complexes containing halogenated molecules | 40 | 4.90 | 129 | 130 |
X40x5 | PECs for X40 | 200 | 4.02 | 129 | 130 |
XB8 | Binding energies of small halogen-bonded dimers | 8 | 5.23 | 131 | |
XB20 | Binding energies of large halogen-bonded dimers | 20 | 6.06 | 131 | |
@@
AE11 | Absolute atomic energies of closed shell atoms from Ar to Rn | 11 | 6.5e6 | 132 | |
AE18 | Absolute atomic energies of hydrogen through argon | 18 | 1.5e5 | 133 | |
AL2X6 | Dimerisation energies of AlX3 compounds | 6 | 36.93 | 10 | 134 |
@@
AlkAtom19 | Atomization energies of n=1-8 alkane | 19 | 1829.31 | 70 | |
ALKBDE10 | Dissociation energies in group-1 and -2 diatomics | 10 | 105.53 | 135 | |
AlkIsod14 | Isodesmic reaction energies in n = 3–8 alkanes. | 14 | 10.35 | 70 | |
BDE99MR | Multi-reference (MR) bond dissociation energies from W4-11 | 16 | 54.51 | 82 | |
BDE99nonMR | Single-reference (SR) bond dissociation energies from W4-11 | 83 | 114.98 | 82 | |
@@
BSR36 | Hydrocarbon bond separation reaction energies | 36 | 19.45 | 10 | 136 |
CR20 | Cycloreversion reaction energies | 20 | 22.31 | 137 | |
DARC | Reaction energies of Diels-Alder reactions | 14 | 34.94 | 10 | 134 |
@@
EA50 | Vertical electron affinities | 50 | 45.42 | 138 | |
FH51 | Reaction energies in various (in-)organic systems | 51 | 45.35 | 139 | 140 |
G21EA | Adiabatic electron affinities | 25 | 41.05 | 141, 142 | 143 |
G21IP | Adiabatic ionization potentials | 36 | 265.65 | 141, 142 | 143 |
G2RC24 | Reaction energies of systems from G2/97 set | 24 | 71.05 | 10 | 143 |
HAT707MR | Heavy-atom transfer energies (MR) from W4-11 | 202 | 83.41 | 82 | |
HAT707nonMR | Heavy-atom transfer energies (SR) from W4-11 | 505 | 74.79 | 82 | |
@@
HNBrBDE18 | Homolytic N-Br bond dissociation energies | 18 | 56.95 | 144 | |
IP23 | Vertical ionization potentials | 23 | 307.03 | 145 | |
IP30 | Vertical ionization potentials | 30 | 281.77 | 146 | |
MB08-165 | Decomposition energies of 165 artificially constructed molecules, each containing 8 atoms | 165 | 154.46 | 147 | |
@@
MX34 | Electronic atomization energies for a range of ionic clusters with the NaCl structure | 34 | 833.23 | 148 | |
@@
P34 | The atomization energies, ionization potentials, and electron affinities of systems containing heavy p-block elements up to xenon | 68 | 508.34 | 149 | |
@@
PlatonicHD6 | Reaction energies of homodesmotic reactions involving platonic hydrocarbon cages, CnHn (n = 4, 6, 8, 10, 12, 20) | 6 | 136.71 | 150 | |
PlatonicID6 | Reaction energies of isodesmic reactions involving platonic hydrocarbon cages, CnHn (n = 4, 6, 8, 10, 12, 20) | 6 | 96.19 | 150 | |
PlatonicIG6 | Reaction energies of isogyric reactions involving platonic hydrocarbon cages, CnHn (n = 4, 6, 8, 10, 12, 20) | 6 | 356.33 | 150 | |
PlatonicTAE6 | Total atomization energies of platonic hydrocarbon cages, CnHn (n = 4, 6, 8, 10, 12, 20) | 6 | 2539.27 | 150 | |
RC21 | Fragmentations and rearrangements in organic radical cations | 21 | 43.94 | 10 | 151 |
RSE43 | Radical stabilization energies | 43 | 9.96 | 10 | 152 |
@@
SN13 | Nucleophilic substitution energies | 13 | 25.67 | 82 | |
@@
WCPT6 | Tautomerization energies for water-catalyzed proton-transfer reactions | 6 | 7.53 | 64 | |
YBDE18 | Bond-dissociation energies in ylides | 18 | 53.11 | 10 | 153 |
@@
3d4dIPSS | Spin-state relative energies and ionization potentials for first-row (3d) and second-row (4d) transition metals | 35 | 132.89 | 154, 155, 156, 157 | |
CUAGAU83 | Atomization, ionization, isomerization, and binding energies of molecular systems containing Cu, Ag, and Au | 83 | 114.12 | 45, 158 | |
DAPD | Difficult bonding energies of Pd-containing diatomic molecules | 12 | 64.93 | 46 | |
MME52 | Reaction energies and barrier heights of metalloenzyme models | 52 | 21.71 | 159 | 35 |
MOBH28 | Barrier heights involving organometallic complexes | 56 | 26.91 | 160 | 43 |
@@
TMD10 | Statistically significant subsets from TMD60, which contain bonding energies for diatomic transition-metal species | 10 | 88.94 | 161 | 162 |
MOR13 | Statistically significant subsets from MOR41, which contain the reaction energies of closed-shell organometallic reactions | 13 | 33.90 | 163 | 162 |
TMB11 | Statistically significant subsets from TMB50, which contain reaction barriers involving transition metal complexes | 11 | 16.14 | 164, 165, 166, 167 | 162 |
End of Table |

### v2

@@ Abstract
We present GSCDB137, a rigorously curated benchmark library of 137 data sets (8377 entries) covering main-group and transition-metal reaction energies and barrier heights, (intramolecular) non-covalent interactions, dipole moments, polarizabilities, electric-field response energies, and vibrational frequencies. Legacy data from GMTKN55 and MGCDB84 have been updated to today’s best reference values; redundant, spin-contaminated, or low-quality points were removed, and many new, property-focused sets were added. Testing 29 popular density functional approximations (DFAs) confirms the expected Jacob’s-ladder hierarchy overall but also reveals notable exceptions: functional performance for frequencies and electric-field properties correlates poorly with that for other ground-state energetics. B97M-V and B97X-V are the most balanced hybrid meta-GGA and hybrid GGA, respectively; B97M-V and revPBE-D4 lead the meta-GGA and GGA classes. Double hybrids lower mean errors by about 25 % versus the best hybrids but demand careful frozen-core, basis-set, and multi-reference treatment. GSCDB137 offers a comprehensive, openly documented platform for stringent DFA validation and for training the next generation of non-empirical and machine-learned functionals.
@@ 1 Introduction
As more research groups began generating high-quality (and sometimes low-quality) benchmark data, the need to systematically integrate and curate these efforts into a larger, validated, and community-accessible database became increasingly apparent. A landmark effort was the GMTKN24 and GMTKN30 databases 37, 38, which cover general main-group thermochemistry, kinetics, and noncovalent interactions. In 2017, the larger GMTKN55 compilation was introduced 10 and is now employed to evaluate nearly all modern density functionals 10, 39, 40. That same year, the even larger Main Group Chemistry Database (MGCDB84) was also presented, providing a comprehensive benchmark of over 200 density functionals 16. Other significant efforts were also reported.41, 42
@@
Given that these state-of-the-art databases are now about eight years old, there is a significant opportunity to improve diversity and quality of data in new compilations. This work reports the result of our effort to develop a larger and more comprehensive database, Gold-Standard Chemical Database 137 (GSCDB137). GSCDB137 aims to deliver gold-standard accuracy and contains 137 datasets encompassing a total of 8377 individual data points (requiring around 14k single-point energy calculations). Key improvements in diversity include extensive transition-metal data drawn from realistic organometallic reactions 43, 44, 45 and well-defined model complexes 46, 47. A second important addition is energy differences that reflect molecular properties. Recently, there has been controversy regarding the accuracy of electron densities from modern functionals 48 and thus we add density-dependent properties, such as dipole moments 49, polarizabilities 50, 51, and the field-dependence of energies 52. Additionally, we introduce datasets for vibrational frequencies 53, further extending the range of available benchmarking targets. Beyond these brand-new energy categories, we also expand the coverage and diversity within established categories. Furthermore, we carefully and systematically prune all questionable data points, especially those potentially affected by spin-symmetry-breaking 54, 55, 56, yielding a benchmark suite aimed at gold-standard accuracy across an unprecedented range of chemistry.
@@
The outline of the paper is as follows. In the next section, we provide an overview of the Gold-Standard Chemical Database 137 (GSCDB137), including a comprehensive description of all 137 benchmark sets, their characteristics, and reference values (Table LABEL:chap5tab:overview). In Section 3, we discuss important considerations in constructing the database and choosing reference values. This includes the integration of existing benchmark databases (MGCDB84 and GMTKN55), updates and adjustments to reference values, resolution of spin contamination issues, inclusion of newly added data sets—particularly for transition metals and molecular properties—and basis set considerations. In Section 4, we present a DFT benchmark study that evaluates the performance of 29 selected density functionals across the new database, identifying the strengths and limitations of the functionals themselves and their corresponding functional design. Finally, in Section 5, we present our conclusions, summarizing the key findings and discussing the implications for future density functional development.
@@ 2 Overview of the Database
Figure 1 summarizes the generation and composition of GSCDB137. Table LABEL:chap5tab:overview is a simple description of all 137 benchmark sets, including a short description, the number of data points contained in each set, the root-mean-square value of their energy differences, and the reference. A more detailed version is attached in the Supplemental Information.
@@
| BH28 | Highly accurate subset chosen from BHPERI, CRBH20, BHDIV10, and PX13 sets | 28 | 35.18 | 57 | |
@@
| BH876 | Comprehensive reaction barrier heights | 876 | 26.93 | 58 | |
@@
| BHPERI11 | Barrier heights of pericyclic reactions | 11 | 23.00 | 59 | 37 |
@@
| CRBH14 | Barrier heights for cycloreversion of heterocyclic rings | 14 | 45.85 | 60 | |
| DBH22 | Highly accurate subset of BH76 | 22 | 29.22 | 61 | 32 |
| INV23 | Inversion/Racemization barrier heights | 23 | 36.57 | 62 | |
| ORBH35 | Difficult barrier heights for oxygen reactions (e.g., vinylperoxy radical and CO3) | 35 | 32.86 | 63 | |
| PX9 | Proton-exchange barriers in H2O, NH3, and HF clusters | 9 | 38.87 | 64 | |
| WCPT26 | Barrier heights of water-catalyzed proton-transfer reactions | 26 | 36.11 | 65 | |
@@
| Dip146 | Dipole moments for 152 small systems | 190 | 0.12 | 49 | |
| HR46 | Static polarizabilities for 46 systems | 128 | 0.47 | 51 | 66 |
| OEEF | Relative energies in oriented external electric fields compared to zero field | 128 | 18.07 | 52 | |
| Pol130 | Static polarizabilities for 132 small systems | 296 | 1.64 | 50, 67 | |
| T144 | Static polarizabilities for 144 systems | 421 | 0.69 | 51 | 68 |
| V30 | Frequencies of small molecular dimers with different polarity combinations (polar-polar, polar-nonpolar, and nonpolar-nonpolar) | 275 | 6.0e-4 | 53 | |
@@
| AlkIsomer11 | Isomerization energies of alkanes with 4-8 carbon atoms | 11 | 1.81 | 69 | |
| C20C246 | Isomerization energies of C20 and C24 | 6 | 41.54 | 70 | |
| C60ISO7 | Isomerization energies of C60 isomers | 7 | 102.22 | 71 | |
| DIE60 | Isomerization energies for double-bond migration in conjugated dienes | 60 | 5.06 | 72 | |
| EIE22 | Isomerization energies of enecarbonyls | 22 | 4.97 | 73 | |
| ISO34 | Isomerisation energies of small- and medium-sized organic molecules | 34 | 20.30 | 10 | 74 |
| ISOL23 | Isomerisation energies of large organic molecules | 23 | 28.03 | 75 | 76 |
| ISOMERI- ZATION20 | Isomerization energies from the W4-11 data set | 20 | 44.05 | 77 | |
| PArel | Isomerization energies of protonated isomers | 20 | 6.23 | 10 | |
| Styrene42 | Isomerization energies of C8H8 | 42 | 68.86 | 78 | |
| TAUT15 | Isomerization energies of tautomers | 15 | 4.52 | 10 | |
| Intramolecular Noncovalent Interactions | |||||
| ACONF | Relative energies in conformers of alkane conformers | 15 | 2.23 | 79 | |
| Amino20x4 | Relative energies in conformers of amino acid conformers | 80 | 2.98 | 80 | |
| BUT14DIOL | Relative energies in conformers of butane-1,4-diol conformers | 64 | 2.91 | 10 | 81 |
| ICONF | Relative energies in conformers of inorganic systems | 17 | 4.54 | 10 | |
@@
| MCONF | Relative energies in conformers of melatonin conformers | 51 | 5.36 | 10 | 82 |
| PCONF21 | Relative energies in conformers of tri- and tetrapeptide conformers | 18 | 1.79 | 10 | 83, 84 |
| Pentane13 | Relative energies in conformers of stationary points on the n-pentane torsional surface | 13 | 6.75 | 85 | |
| SCONF | Relative energies in conformers of sugar conformers | 17 | 4.96 | 10 | 86 |
| UPU23 | Relative energies in conformers of RNA-backbone conformers | 23 | 6.52 | 87 | |
@@
| 3B-69 | 3-Body non-additive interaction energies in three different orientations of 23 molecular crystals | 69 | 0.69 | 88 | |
| 3BHET | 3-Body non-additive interaction energies in molecular trimers | 20 | 2.01 | 89 | |
| A19Rel6 | Relative energies of 19 complexes from the A24 set on potential energy curves (PECs) | 114 | 1.65 | 90 | |
| A24 | Highly accurate binding energies of small non-covalent complexes | 24 | 2.65 | 91 | 92 |
| ADIM6 | Interaction energies of n-alkane dimers | 6 | 3.66 | 10 | 93 |
| AHB21 | Interaction energies in anion-neutral dimers | 21 | 26.17 | 94 | |
| Bauza30 | Binding energies of halogen-, chalcogen-, and pnicogen-bonded dimers | 30 | 23.65 | 95 | 96 |
| BzDC215 | PECs for benzene interacting with two rare-gas atoms and eight first- and second-row hydrides | 215 | 1.81 | 97 | |
@@
| CHB6 | Interaction energies in cation-neutral dimers | 6 | 27.85 | 94 | |
| CT20 | Binding energies of charge-transfer complexes | 20 | 1.07 | 98 | |
| DS14 | Binding energies of complexes containing divalent sulfur | 14 | 3.70 | 99 | |
| FmH2O10 | Binding energies of isomers of F-(H2O)10 | 10 | 168.50 | 100 | 101 |
| H2O16Rel4 | Relative energies in conformers of (H2O)16 (boat and fused cube structures) | 4 | 0.45 | 102 | |
| H2O20Rel9 | Relative energies in conformers of (H2O)20 (low-energy structures) | 9 | 2.76 | 100 | 103 |
@@
| HEAVY28 | Binding energies between heavy element hydrides | 28 | 1.43 | 10 | 93 |
@@
| HW6Cl5 | Binding energies of Cl-(H2O)n (n = 1-6) | 5 | 62.84 | 100 | 101 |
| HW6F | Binding energies of F-(H2O)n (n =1-6) | 6 | 81.42 | 100 | 101 |
@@
| IL16 | Interaction energies in anion-cation dimers | 16 | 109.34 | 94 | |
@@
| S66Rel7 | Relative energies of 66 complexes from the S66 set on PECs | 462 | 2.74 | 124 | 35 |
| Shields38 | Binding energies of (H2O)n (n = 2 - 10) | 38 | 51.61 | 125, 126 | 127 |
| SW49Bind22 | Binding energies of isomers of SO42-(H2O) (n = 3 -5) | 22 | 34.20 | 128 | |
| SW49Rel28 | Relative energies in conformers of SO42-(H2O)n (n = 3-5) | 28 | 1.55 | 128 | |
| TA13 | Binding energies of dimers involving radicals | 13 | 22.00 | 129 | |
| WATER27 | Binding energies in (H2O)n, H+(H2O)n and OH-(H2O)n | 27 | 101.19 | 125, 126 | 130 |
| X40 | Binding energies of complexes containing halogenated molecules | 40 | 4.90 | 131 | 132 |
| X40x5 | PECs for X40 | 200 | 4.02 | 131 | 132 |
| XB25 | Binding energies of halogen-bonded dimers | 25 | 5.88 | 133 | |
@@
| AE11 | Absolute atomic energies of closed shell atoms from Ar to Rn | 11 | 6.5e6 | 134 | |
| AE18 | Absolute atomic energies of hydrogen through argon | 18 | 1.5e5 | 135 | |
| AL2X6 | Dimerisation energies of AlX3 compounds | 6 | 36.93 | 10 | 136 |
@@
| AlkAtom19 | Atomization energies of n=1-8 alkane | 19 | 1829.31 | 69 | |
| ALKBDE10 | Dissociation energies in group-1 and -2 diatomics | 10 | 105.53 | 137 | |
| AlkIsod14 | Isodesmic reaction energies in n = 3–8 alkanes. | 14 | 10.35 | 69 | |
| BDE99MR | Multi-reference (MR) bond dissociation energies from W4-11 | 16 | 54.51 | 77 | |
| BDE99nonMR | Single-reference (SR) bond dissociation energies from W4-11 | 83 | 114.98 | 77 | |
@@
| BSR36 | Hydrocarbon bond separation reaction energies | 36 | 19.45 | 10 | 138 |
| CR20 | Cycloreversion reaction energies | 20 | 22.31 | 139 | |
| DARC | Reaction energies of Diels-Alder reactions | 14 | 34.94 | 10 | 136 |
@@
| EA50 | Vertical electron affinities | 50 | 45.42 | 140 | |
| FH51 | Reaction energies in various (in-)organic systems | 51 | 45.35 | 141 | 142 |
| G21EA | Adiabatic electron affinities | 25 | 41.05 | 143, 144 | 145 |
| G21IP | Adiabatic ionization potentials | 36 | 265.65 | 143, 144 | 145 |
| G2RC24 | Reaction energies of systems from G2/97 set | 24 | 71.05 | 10 | 145 |
| HAT707MR | Heavy-atom transfer energies (MR) from W4-11 | 202 | 83.41 | 77 | |
| HAT707nonMR | Heavy-atom transfer energies (SR) from W4-11 | 505 | 74.79 | 77 | |
@@
| HNBrBDE18 | Homolytic N-Br bond dissociation energies | 18 | 56.95 | 146 | |
| IP23 | Vertical ionization potentials | 23 | 307.03 | 147 | |
| IP30 | Vertical ionization potentials | 30 | 281.77 | 148 | |
| MB08-165 | Decomposition energies of 165 artificially constructed molecules, each containing 8 atoms | 165 | 154.46 | 149 | |
@@
| MX34 | Electronic atomization energies for a range of ionic clusters with the NaCl structure | 34 | 833.23 | 150 | |
@@
| P34AE | The atomization energies of systems containing heavy p-block elements up to xenon | 44 | 617.89 | 151 | |
| P34EA | The electron affinities of systems containing heavy p-block elements up to xenon | 9 | 44.28 | 151 | |
| P34IP | The ionization potentials of systems containing heavy p-block elements up to xenon | 15 | 224.49 | 151 | |
@@
| PlatonicRE18 | Reaction energies of homodesmotic, isodesmic, and isogyric reactions involving platonic hydrocarbon cages, CnHn (n = 4, 6, 8, 10, 12, 20) | 18 | 227.24 | 152 | |
| PlatonicTAE6 | Total atomization energies of platonic hydrocarbon cages, CnHn (n = 4, 6, 8, 10, 12, 20) | 6 | 2539.27 | 152 | |
| RC21 | Fragmentations and rearrangements in organic radical cations | 21 | 43.94 | 10 | 153 |
| RSE43 | Radical stabilization energies | 43 | 9.96 | 10 | 154 |
@@
| SN13 | Nucleophilic substitution energies | 13 | 25.67 | 77 | |
@@
| WCPT6 | Tautomerization energies for water-catalyzed proton-transfer reactions | 6 | 7.53 | 65 | |
| YBDE18 | Bond-dissociation energies in ylides | 18 | 53.11 | 10 | 155 |
@@
| 3d4dIPSS | Spin-state relative energies and ionization potentials for first-row (3d) and second-row (4d) transition metals | 32 | 138.97 | 156, 157, 158, 159 | |
| CUAGAU83 | Atomization, ionization, isomerization, and binding energies of molecular systems containing Cu, Ag, and Au | 83 | 114.12 | 46, 160 | |
| DAPD | Difficult bonding energies of Pd-containing diatomic molecules | 12 | 64.93 | 47 | |
| MME52 | Reaction energies and barrier heights of metalloenzyme models | 52 | 21.71 | 161 | 35 |
| MOBH28 | Barrier heights involving organometallic complexes compiled from real-life organometallic catalytic problems | 56 | 26.91 | 162 | 43 |
@@
| TMD10 | Statistically significant subsets from TMD60, which contain bonding energies for diatomic transition-metal species | 10 | 88.94 | 163 | 164 |
| MOR13 | Statistically significant subsets from MOR41, which contain the reaction energies of closed-shell organometallic reactions | 13 | 33.90 | 45 | 164 |
| TMB11 | Statistically significant subsets from TMB50, which contain reaction barriers involving transition metal complexes | 11 | 16.14 | 165, 166, 167, 168 | 164 |
| End of Table | |||||